Please draw:

-starting materials for stilbene synthesis with the Wittig reaction mechanism; which will be one aldehyde + one ylide compound

-the arrow-pushing mechanism for the ylide formation.

-the arrow-pushing mechanism for the formation of stilbene

Please be specific in the compounds, in other words not with R' groups.

Thank you!
Answer #1

Reagents used for Ylide formation ph ph Ph₂ p ph Triphenyl Phosphate and, Li-(H₂-CH₂-CH2-CH3 Bulis for synthesis of Stilbene,Ylide formation: (Hq Br Ph Ph₃P РЬ — Р CH Buli ph Mechanism: ☺ SN? reaction : HH Ph ph -Ć - Br ph. CH2 CP : Phi Ph 으 5 РЬ + BYlides reacts with aldehydes to give alkenes, called Wittig Reaction. for. synthesis of Stilbene we use, 0 =CH- (BenzaldehydeStep 3) | Reverse (2+2) Cycloaddition P.Ph3 CH CH to=P-Pha Stilbene